Use the formula that I sent you, if you show me the steps clear and the answer is correct, you will get thumbs up.
A $1,000 par value bond with seven years lert to maturity has a 9 percent coupon rate (paid semiannually) and is selling for $945.80. What is its yield to maturity? (An equation i sufficient.) 4.
3. Coupon Bond Using the same strategy used for the fixed-payment loan: P price of coupon bond C-yearly coupon payment F face value of the boned n = years to maturity date 1+i (İH)2 (1+が (1+i) (1+i) 2.
Answer #1

YTM = interest +(redemption or face value-market value)/n / (redemption or face value+market value)/2

45+(1000-945.80)/18 / (1000+945.80)/2

YTM (semi annual) = 41.99/972.9 = .0431*100 = 4.31%

YTM annual = 4.31*2 =8.62%